OpenStudy (anonymous):

explain why the function f(x) = x^3 +2x-4 only has a root

OpenStudy (anonymous):

because it is a depressed cubic

OpenStudy (anonymous):

LOL what does that mean?

OpenStudy (anonymous):

Not sure if I understand the question. There are 3 roots, 1 real, 2 complex

OpenStudy (anonymous):

A depressed cubic was extensively analyzed by Italian mathematicians in the 16th century and their work was picked up by Descartes and euler in the following century. a standard cubic is in the form ax^3+bx^2+cx+d a depressed cubic has no x^2 term in it

OpenStudy (anonymous):

LOL thank you mr historian!

OpenStudy (anonymous):

suppose we have three real solutions. \[x1+x2+x3=0\] \[x1.x2+x2.x3+x3.x2=2 \] so \[(x1+x2+x3)^{2}=x1^{2}+x2^{2}+x3^{2}+2(x1x2+x2x3+x3x1)=0\] so \[x1^{2}+x2^{2}+x3^{2}+2(2)=0\] that's impossible so our soulutions are not real,then just one is real

OpenStudy (anonymous):

suppose that this function has 2 root, we call them that are \[x_1\] and \[x_2\]. Because they are root, we have \[x^3_1+2x_1-4=x^3_2+2x_2-4\] Then we can deduce that \[x_1=x_2\]